Unexposed silver halides are removed from photographic film when they react with sodium thiosulfate (Na2S2O3, called hypo) to form the complex ion Ag(S2O3)23-. The reaction with silver bromide is:

AgBr(s) + 2 S2O32- (aq) ⇌ Ag(S2O3)23- (aq) + Br- (aq)

The equilibria that need to be considered are:

Ag+ + 2 S2O32- (aq) ⇌ Ag(S2O3)23- (aq) Kf = 4.7 x 1013

AgBr(s) ⇌ Ag+ (aq) + Br- (aq) Ksp = 5.0 x 10-13

What mass of Na2S2O3 is required to prepare 1.00 L of a solution that will dissolved 1.00 of AgBr by the formation of Ag(S2O3)23-?
Answer #1

Ag+ + 2 S2O32- (aq) ⇌ Ag(S2O3)23- (aq) Kf = 4.7 x 1013

AgBr(s) ⇌ Ag+ (aq) + Br- (aq) Ksp = 5.0 x 10-13

Combining equation above 2 equations that cancel out silver ions : AgBr(s) + 2 S2O32- (aq) ⇌ Ag(S2O3)23- (aq) + Br- (aq)

Ksp = [Ag+] [Br-] = 5.0 x 10-13

Molar mass of AgBr = 187.77 g/mol

Molarity of AgBr . = [Br-] = wt/molar mass)/vol in lts = (1 g/187.77 g/mol)/1 lt = 0.00533 M

Ksp = [Ag+] [Br-] = 5.0 x 10-13

[Ag+] = 5 x 10-13 / 5.33 x 10-3 = 0.94 x 10-10 M

   Ag+ + 2 S2O32- (aq) ⇌ Ag(S2O3)23- (aq)

At equilibrium , [Ag(S2O3)23- ] = 5.33 x 10-3 as AgBr is dissolved

Kf = 4.7 x 1013 = [Ag(S2O3)23- ]/ [Ag+] [S2O32- ]2 = 5.33 X 10-3/0.94 x 10-10 [S2O32- ]2

[[S2O32- ] = 5.67 x 107/4.7 X 1013 = 1.1 x 10-3 M

Total no. of moles of thiosulphate ion = amount of thiosulphate that reacted with silver ions and free thiosulphate ions =

in [Ag(S2O3)23- ] 2 thiosulphate ions + free thiosulphate ion = 2 x 0.00533 + 0.0011 = 0.01176 M

Volume is 1 litre

So, no. of moles of thiosulphate ions = 0.01176 moles

Molar mass of Na2S2O3 = 158.11 g/mol

Mass of Na2S2O3 required = Moles x molar mass = 0.01176 moles x 158.11 g/mol = 1.86 gms